OpenStudy (anonymous):

assume fixed costs equal $300,000. The price per unit equals $10 and the variable cost per unit equals $7. Find the breakeven point in terms of both units sold and sales revenue

OpenStudy (gw2011):

Let S=Sales Let FC=Fixed Costs Let VC=Variable Costs Let x=Units Then, S=FC+VC S=10x FC=$300,000 VC=3x 10x=300,000+7x 3x=300,000 x=100,000 units (10)(100,000)=300,000+(7)(100,000) $1,000,000=$1,000,000 The breakeven point is 100,000 units and $1,000,000 because the 100,000 units make the Sales=The Total Costs